An unstable high-energy particle is created in the laboratory, and it moves at a speed of 0.989c. Relative to a stationary reference frame fixed to the laboratory, the particle travels a distance of 1.77 x 10 m before disintegrating. What is (a) the proper distance and (b) the distance measured by a hypothetical person traveling with the particle? Determine the particles (c) proper lifetime and (d) its dilated lifetime (a) Number Units (b) Number Units (c) Number nits (d) Number Units
Answer #1

(a)

The proper distance is L = 1.77X10-3 m

***************************************************************************************************

(b)

The distance measured by a hypothetical person traveling with the particle will be,

L' = L√[1-v2/c2] = (1.77X10-3 m)√[1-(0.989c)2/c2]

or, L' = 2.62X10-4 m

***************************************************************************************************

(c)

The proper lifetime will the time measured by the hypothetical person for the particle to disintegrate...

t = L'/v = (2.62X10-4 m)/(0.989X3X108 m/s)

or, t = 8.82X10-13s is the proper lifetime.

***************************************************************************************************

(d)

Dilatedl lifetime is,

t' = t/√[1-(0.989c)2/c2] = (8.82X10-13s)/√[1-(0.989c)2/c2]

or, t' = 5.96X10-12s

***************************************************************************************************
This concludes the answers. Check the answer and let me know if it's correct. If you need any more clarification, modification or correction, feel free to ask.....